# CPathAgent: An Agent-based Foundation Model for Interpretable High-Resolution Pathology Image Analysis Mimicking Pathologists’ Diagnostic Logic

### v1

@@ B.2 Comparison of WSI Classification Performance Among Pathology-specific LMMs
Pathology-specific Large Multimodal Models
Quilt-LLaVA 55.3 54.0 53.3 58.6 61.9 45.8 54.8
PathGen-LLaVA 66.5 53.5 59.0 78.6 54.7 49.2 60.2
@@
[…]
B.3 Ablation Study on Applying Gemini-2.5-Pro as an Agent for WSI Tasks
@@
Results: Interestingly, while this approach yields modest improvements on TCGA-BLCA and TCGA-THCA datasets, the overall performance decreases by 6.7% (from 72.1% to 65.4%), as shown in the "Agent-based Approach" part of Table 4. This contrasts with the findings in Section 4.2 of the main paper, where applying agent-based approaches to advanced LMMs on large regions VQA demonstrated performance gains.
@@
This finding validates our progressive training data generation framework, which leverages comprehensive WSI reports to guide region-specific descriptions. These targeted descriptions enable us to prompt Gemini-2.5-Pro to capture multi-scale morphological features across different viewing perspectives. By understanding which anatomical locations contain specific pathological features, we can further prompt Gemini-2.5-Pro to generate higher-quality navigation paths with enhanced spatial awareness, which in turn enables more precise and accurate multi-scale, multi-view reasoning generation. This creates a hierarchical sequence where each stage builds upon the quality of previous stages, resulting in cascading improvements throughout the entire data generation pipeline.
@@ C.1.3 Details for WSI Task-Specific Fine-Tuning
Model Architecture. The MIL framework commonly used for WSI classification includes three learnable components: (1) a fully-connected layer to reduce the dimensionality of features to 256, (2) an attention network to aggregate and transform the instance features, and (3) a final fully-connected layer for making predictions. We experiment with ABMIL [61] and DSMIL [5], both of which share the same fully-connected layers for reducing feature dimensionality and making predictions. For the attention network, ABMIL uses a gated attention mechanism, while DSMIL introduces a dual-stream architecture.
C.2 Hardware and Training Cost of CPathAgent
@@ C.3 Training Hyperparameters
Appendix D Examples and Prompts
D.1 Examples of Region Descriptions Extracted from WSI Report Pairs
@@
As shown in Figure 14, we extract relevant region descriptions from WSI reports by identifying specific diagnostic findings in targeted areas. This process systematically parses pathological reports to highlight key diagnostic information related to the given region, addressing the limitation in TCGA datasets where WSI reports exist but region-specific descriptions are missing.
D.2 Examples of PathMMU-HR² Samples
@@
Figure 15 shows the examples of PathMMU-HR² dataset. These VQA samples demonstrate exceptional quality through the requirement for multi-scale, multi-view reasoning to arrive at accurate diagnoses. As illustrated in the examples, the correct identification of conditions such as fibromatosis (desmoid tumor) and well-demarcated clear cell tumors necessitates careful examination of architectural patterns and cellular features across different magnification levels. The diagnostic process requires the model to analyze tissue structure at lower magnifications while progressively zooming in to evaluate the cellular morphology and stromal characteristics at higher magnifications. This multi-scale analytical approach reflects real-world pathological practice, where definitive diagnoses depend on the integration of macro-architectural features with microscopic cellular details.
D.3 Examples of Region Selection Process Based on WSI Overview
@@
Figure 16 and Figure17 demonstrate the generated region selection results. As illustrated, the model segments the WSI overview into distinct regions and identifies which areas require high-magnification examination. For example, in Figure 16, the model correctly prioritizes core tumor areas for detailed analysis while also identifying critical tumor interface/periphery regions and margin assessment areas to evaluate potential residual tumor presence. In addition, the model appropriately excludes the benign/background breast tissue areas from high-magnification requirements, as these regions primarily consist of adipose tissue and fibrous stroma that appear benign and uninvolved at low magnification, thus optimizing the diagnostic workflow by focusing computational resources on clinically significant regions.
@@
D.4 Examples of Navigation Paths and Multi-Scale Reasoning Generated by CPathAgent
D.4.1 General Navigation Path Planning and Reasoning
@@
Figures 18 and 19 show how CPathAgent navigates and analyzes pathology images like an expert pathologist. For example, as illustrated in Figure 18, CPathAgent begins with a comprehensive overview at low magnification (1.0x), scanning the entire tissue to identify distinct regions and architectural patterns. In this case, it recognizes adipose tissue on the left and dense cellular areas on the right, establishing the overall tissue landscape.
@@
D.4.2 VQA-oriented Navigation Path Planning and Reasoning
Figure 20 demonstrates CPathAgent’s question-guided reasoning approach for VQA task. When presented with a specific question about "morphological patterns describing the relationship between neoplastic cell populations and stromal elements at their interface," CPathAgent tailors its entire analysis strategy to address this inquiry.
@@ D.4.2 VQA-oriented Navigation Path Planning and Reasoning
Figure 14: Examples of region description extraction from WSI reports.
Figure 15: Examples of VQAs from the PathMMU-HR² dataset.
Figure 16: An example of region selection based on a WSI overview.
Figure 17: An example of region selection based on a WSI overview.
Figure 18: An example of CPathAgent’s navigation path planning and multi-scale, multi-view reasoning process.
Figure 19: An example of CPathAgent’s navigation path planning and multi-scale, multi-view reasoning process.
Figure 20: An example of CPathAgent’s VQA-oriented navigation path planning and multi-scale, multi-view reasoning process.
D.5 Prompts
@@
Gemini-2.5-Pro Prompts for Dataset Construction: We design a series of prompts to guide Gemini-2.5-Pro in constructing our pathologist-like planning and reasoning dataset generation pipeline. Figure 21 presents the prompt used to extract region-relevant descriptions from WSI reports, while Figure 22 shows the prompt for region selection using WSI overview images and paired reports. For multi-scale image description generation, Figure 23 displays the prompt that generates detailed descriptions for image patches at different magnifications. Figure 24 contains the prompt for creating navigation plans using coordinate-annotated region overviews and multi-scale descriptions. Finally, Figure 25 shows the prompt that generates pathologist-like diagnostic reasoning based on extracted patches from navigation paths and paired region descriptions.
@@
For VQA dataset construction, we use several specialized prompts: Figure 26 shows the prompt that instructs the model to generate VQA samples, Figure 27 presents the prompt for predicting question answers from text-only inputs (applied to both Gemini-2.5-Pro and DeepSeek-V3), Figure 28 displays the prompt that guides the model to create question-guided navigation plans, and Figure 29 contains the prompt for generating reasoning-based answers.
@@
CPathAgent Training and Inference Prompts: The prompts for CPathAgent follow a similar design, while removing the additional information utilized by Gemini-2.5-Pro (such as pathology reports), focusing on vision-only pathological analysis that replicates clinical diagnostic processes. Figures 30, 31, and 32 present the core prompts used for CPathAgent’s training and inference. These prompts guide CPathAgent to perform navigation planning and reasoning using only visual input. The VQA-specific prompts shown in Figures 33 and 34 display the prompts that used for CPathAgent’s question-guided navigation and reasoning.
@@
Figure 21: Prompt for Gemini-2.5-Pro to extract description related to a given region from the corresponding WSI paired pathology report.
@@
Figure 22: Prompt for Gemini-2.5-Pro to generate region selection results given a WSI overview and its paired pathology report.
@@
Figure 23: Prompt for Gemini-2.5-Pro to generate descriptions for 1 image at 1X scale, 4 images at 2X scale, and 16 images at 4X scale, given these images and their paired region descriptions.
@@
Figure 24: Prompt for Gemini-2.5-Pro to generate navigation path planning given the region overview with coordinates and 21 multi-scale image descriptions generated in the previous step.
@@
Figure 25: Prompt for Gemini-2.5-Pro to generate multi-scale multi-image reasoning for description and diagnosis based on the planned navigation path, given the cropped images along the navigation path and the overall region description extracted from the previous step.
@@
Figure 26: Prompt for Gemini-2.5-Pro to generate VQA samples given the region overview and the 21 multi-scale image descriptions generated from the previous step.
@@
Figure 27: Prompt for Gemini-2.5-Pro to conduct educated guesses given only the text portion of the generated VQA samples.
@@
Figure 28: Prompt for Gemini-2.5-Pro to generate navigation path planning given the region overview with coordinates, 21 multi-scale image descriptions and the question generated in the previous step.
@@
Figure 29: Prompt for Gemini-2.5-Pro to generate VQA-oriented reasoning based on the planned navigation path, given the cropped images along the navigation path, the overall region description extracted from the previous step, and the target question.
@@
Figure 30: Prompt for CPathAgent to generate region selection results given a WSI overview.
@@
Figure 31: Prompt for CPathAgent to generate navigation path planning given the region overview with coordinates.
@@
Figure 32: Prompt for CPathAgent to generate multi-scale multi-image reasoning for description and diagnosis based on the planned navigation path, given the cropped images along the navigation path.
@@
Figure 33: Prompt for CPathAgent to generate VQA-oriented navigation path planning given the region overview with coordinates and the question.
@@
Figure 34: Prompt for CPathAgent to generate VQA-oriented reasoning based on the planned navigation path, given the cropped images along the navigation path and the target question.

### v2

@@ B.4 Ablation Study on Agent Components
B.5 Ablation Study on Applying Gemini-2.5-Pro as an Agent for WSI Tasks
@@
Results: Interestingly, while this approach yields modest improvements on TCGA-BLCA and TCGA-THCA datasets, the overall performance decreases by 6.7% (from 72.1% to 65.4%), as shown in the "Agent-based Approach" part of Table A6. This contrasts with the findings in Section 4.2 of the main paper, where applying agent-based approaches to advanced LMMs on large regions VQA demonstrated performance gains.
@@
This finding validates our progressive training data generation framework, which leverages comprehensive WSI reports to guide region-specific descriptions. These targeted descriptions enable us to prompt Gemini-2.5-Pro to capture multi-scale morphological features across different viewing perspectives. By understanding which locations contain specific pathological features, we can further prompt Gemini-2.5-Pro to generate higher-quality navigation paths with enhanced spatial awareness, which in turn enables more precise and accurate multi-scale, multi-view reasoning generation. This creates a hierarchical sequence where each stage builds upon the quality of previous stages, resulting in cascading improvements throughout the entire data generation pipeline.
@@ C.1.3 Details for WSI Task-Specific Fine-Tuning
Model Architecture. The MIL framework commonly used for WSI classification includes three learnable components: (1) a fully-connected layer to reduce the dimensionality of features to 256, (2) an attention network to aggregate and transform the instance features, and (3) a final fully-connected layer for making predictions. We experiment with ABMIL [4] and DSMIL [5], both of which use the same fully-connected layers for reducing feature dimensionality and making predictions. For the attention network, ABMIL uses a gated attention mechanism, while DSMIL introduces a dual-stream architecture.
C.2 Details for Patch Level Benchmark
PathMMU [48]: An expert-validated pathology benchmark comprising 33,428 multimodal multiple-choice questions and 24,067 images from diverse sources including PubMed, pathology atlases, expert social media posts, and educational content. Constructed using GPT-4V with rigorous validation by seven pathologists, it covers multiple organ systems and pathology subjects for expert-level evaluation of LMMs’ pathology understanding and reasoning capabilities.
PathMMU-Pro [63]: Building upon PathMMU, PathMMU-Pro introduces more strict assessment to address a critical limitation in multimodal evaluation: the potential for models to solve questions using text-only reasoning without truly leveraging visual information. PathMMU-Pro employs a systematic two-step refinement process: (1) training specialized text-only "question-guessing" models that receive only textual inputs to identify and filter out questions answerable without examining pathology images, and (2) enhancing remaining questions by generating more confusing options that are textually similar but require careful visual examination to distinguish, thereby preventing models from guessing answers through superficial text patterns. This enhanced benchmark provides a more strict assessment of models’ intrinsic visual reasoning capabilities in pathology.
C.3 Hardware and Training Cost of CPathAgent
@@ C.4 Training Hyperparameters
Appendix D Limitations
While CPathAgent successfully demonstrates the feasibility of mimicking pathologists’ diagnostic reasoning pathways, several limitations merit future improvement:
1. Synthetic diagnostic pathways. Currently, the diagnostic viewing paths employed by CPathAgent are synthetically generated rather than derived from real pathologist workflows. Although our synthetic paths are designed based on established diagnostic principles and demonstrate effectiveness in our experiments, they may not fully capture the nuanced decision-making processes and adaptive strategies employed by experienced pathologists in clinical practice. Future work should incorporate authentic diagnostic paths collected from pathologists through eye-tracking studies or explicit annotation of their reasoning sequences.
2. Limited task scope. CPathAgent currently focuses exclusively on diagnostic classification tasks. While this represents a critical clinical application, computational pathology encompasses a broader spectrum of tasks including prognostic prediction and biomarker detection. Extending CPathAgent to these domains is theoretically feasible and would significantly expand its clinical utility.
3. Training data scale. Our model was trained on 5,254 WSIs, which, while sufficient to demonstrate proof-of-concept, remains considerably smaller than datasets used by state-of-the-art foundation models such as PRISM (587,196 WSIs) and TITAN (335,645 WSIs). This disparity in scale may limit the model’s generalization capabilities and its ability to capture rare morphological patterns. Scaling up the training data represents a clear path toward improved performance and robustness.
Appendix E Examples and Prompts
E.1 Examples of Region Descriptions Extracted from WSI Report Pairs
@@
As shown in Figure A10, we extract relevant region descriptions from WSI reports by identifying specific diagnostic findings in targeted areas. This process systematically parses pathological reports to highlight key diagnostic information related to the given region, addressing the limitation in TCGA datasets where WSI reports exist but region-specific descriptions are missing.
E.2 Examples of PathMMU-HR² Samples
@@
Figure A11 shows the examples of PathMMU-HR² dataset. These VQA samples require multi-scale, multi-view reasoning to arrive at accurate diagnoses. As illustrated in the examples, the correct identification of conditions such as fibromatosis (desmoid tumor) and well-demarcated clear cell tumors necessitates careful examination of architectural patterns and cellular features across different magnification levels. The diagnostic process requires the model to analyze tissue structure at lower magnifications while progressively zooming in to evaluate the cellular morphology and stromal characteristics at higher magnifications. This multi-scale analytical approach reflects real-world pathological practice, where definitive diagnoses depend on the integration of macro-architectural features with microscopic cellular details.
E.3 Examples of Region Selection Process Based on WSI Overview
@@
Figure A12 and FigureA13 demonstrate the generated region selection results. As illustrated, the model segments the WSI overview into distinct regions and identifies which areas require high-magnification examination. For example, in Figure A12, the model correctly prioritizes core tumor areas for detailed analysis while also identifying critical tumor interface/periphery regions and margin assessment areas to evaluate potential residual tumor presence. In addition, the model appropriately excludes the benign/background breast tissue areas from high-magnification requirements, as these regions primarily consist of adipose tissue and fibrous stroma that appear benign and uninvolved at low magnification, thus optimizing the diagnostic workflow by focusing computational resources on clinically significant regions.
@@
E.4 Examples of Navigation Paths and Multi-Scale Reasoning Generated by CPathAgent
E.4.1 General Navigation Path Planning and Reasoning
@@
Figures A14 and A15 show how CPathAgent navigates and analyzes pathology images like an expert pathologist. For example, as illustrated in Figure A14, CPathAgent begins with a comprehensive overview at low magnification (1.0x), scanning the entire tissue to identify distinct regions and architectural patterns. In this case, it recognizes adipose tissue on the left and dense cellular areas on the right, establishing the overall tissue landscape.
@@
E.4.2 VQA-oriented Navigation Path Planning and Reasoning
Figure A16 demonstrates CPathAgent’s question-guided reasoning approach for VQA task. When presented with a specific question about "morphological patterns describing the relationship between neoplastic cell populations and stromal elements at their interface," CPathAgent tailors its entire analysis strategy to address this inquiry.
@@ E.4.2 VQA-oriented Navigation Path Planning and Reasoning
Figure A10: Examples of region description extraction from WSI reports.
Figure A11: Examples of VQAs from the PathMMU-HR² dataset.
Figure A12: An example of region selection based on a WSI overview.
Figure A13: An example of region selection based on a WSI overview.
Figure A14: An example of CPathAgent’s navigation path planning and multi-scale, multi-view reasoning process.
Figure A15: An example of CPathAgent’s navigation path planning and multi-scale, multi-view reasoning process.
Figure A16: An example of CPathAgent’s VQA-oriented navigation path planning and multi-scale, multi-view reasoning process.
E.5 Prompts
@@
Gemini-2.5-Pro Prompts for Dataset Construction: We design a series of prompts to guide Gemini-2.5-Pro in constructing our pathologist-like planning and reasoning dataset generation pipeline. Figure A17 presents the prompt used to extract region-relevant descriptions from WSI reports, while Figure A18 shows the prompt for region selection using WSI overview images and paired reports. For multi-scale image description generation, Figure A19 displays the prompt that generates detailed descriptions for image patches at different magnifications. Figure A20 contains the prompt for creating navigation plans using coordinate-annotated region overviews and multi-scale descriptions. Finally, Figure A21 shows the prompt that generates pathologist-like diagnostic reasoning based on extracted patches from navigation paths and paired region descriptions.
@@
For VQA dataset construction, we use several specialized prompts: Figure A22 shows the prompt that instructs the model to generate VQA samples, Figure A23 presents the prompt for predicting question answers from text-only inputs (applied to both Gemini-2.5-Pro and DeepSeek-V3), Figure A24 displays the prompt that guides the model to create question-guided navigation plans, and Figure A25 contains the prompt for generating reasoning-based answers.
@@
CPathAgent Training and Inference Prompts: The prompts for CPathAgent follow a similar design, while removing the additional information utilized by Gemini-2.5-Pro (such as pathology reports), focusing on vision-only pathological analysis that replicates clinical diagnostic processes. Figures A26, A27, and A28 present the core prompts used for CPathAgent’s training and inference. These prompts guide CPathAgent to perform navigation planning and reasoning using only visual input. The VQA-specific prompts shown in Figures A29 and A30 display the prompts that used for CPathAgent’s question-guided navigation and reasoning.
@@
Figure A17: Prompt for Gemini-2.5-Pro to extract description related to a given region from the corresponding WSI paired pathology report.
@@
Figure A18: Prompt for Gemini-2.5-Pro to generate region selection results given a WSI overview and its paired pathology report.
@@
Figure A19: Prompt for Gemini-2.5-Pro to generate descriptions for 1 image at 1X scale, 4 images at 2X scale, and 16 images at 4X scale, given these images and their paired region descriptions.
@@
Figure A20: Prompt for Gemini-2.5-Pro to generate navigation path planning given the region overview with coordinates and 21 multi-scale image descriptions generated in the previous step.
@@
Figure A21: Prompt for Gemini-2.5-Pro to generate multi-scale multi-image reasoning for description and diagnosis based on the planned navigation path, given the cropped images along the navigation path and the overall region description extracted from the previous step.
@@
Figure A22: Prompt for Gemini-2.5-Pro to generate VQA samples given the region overview and the 21 multi-scale image descriptions generated from the previous step.
@@
Figure A23: Prompt for Gemini-2.5-Pro to conduct educated guesses given only the text portion of the generated VQA samples.
@@
Figure A24: Prompt for Gemini-2.5-Pro to generate navigation path planning given the region overview with coordinates, 21 multi-scale image descriptions and the question generated in the previous step.
@@
Figure A25: Prompt for Gemini-2.5-Pro to generate VQA-oriented reasoning based on the planned navigation path, given the cropped images along the navigation path, the overall region description extracted from the previous step, and the target question.
@@
Figure A26: Prompt for CPathAgent to generate region selection results given a WSI overview.
@@
Figure A27: Prompt for CPathAgent to generate navigation path planning given the region overview with coordinates.
@@
Figure A28: Prompt for CPathAgent to generate multi-scale multi-image reasoning for description and diagnosis based on the planned navigation path, given the cropped images along the navigation path.
@@
Figure A29: Prompt for CPathAgent to generate VQA-oriented navigation path planning given the region overview with coordinates and the question.
@@
Figure A30: Prompt for CPathAgent to generate VQA-oriented reasoning based on the planned navigation path, given the cropped images along the navigation path and the target question.